# Interactive and Hybrid Imitation Learning: Provably Beating Behavior Cloning

### v2

@@ 3.2 Experimental Comparison
We conduct a simple simulation study comparing the sample efficiency of log-loss Behavior Cloning foster2024behavior and Stagger in four MuJoCo todorov2012mujoco; brockman2016openai continuous control tasks with $H=1000$ and pretrained deterministic MLP experts schulman2015trust; schulman2017proximal. Considering MuJoCo’s low sensitivity to horizon length foster2024behavior, we reveal expert states one by one along consecutive trajectories for BC to allow fine-grained state-wise sample complexity comparison, while Stagger queries exactly one state per iteration by sampling from the latest policy’s rollout and updating immediately with the expert’s annotation. In Stagger, we implement the online learning oracle $\mathbb{A}$ so that it outputs a policy that approximately minimizes the log loss. In addition to log loss, we also include results with online learning oracle minimizing historical examples’ total square loss in Appendix G.2. We defer other implementation details to Appendix G.
@@
Figure 1: State-wise sample complexity comparison between Behavior Cloning and Stagger. Shaded areas show the 10th–90th percentile bootstrap confidence intervals diciccio1996bootstrap over 10 runs. Stagger matches or exceeds BC with $50\%$ fewer annotations, achieving better state-wise annotation efficiency.
@@ 4 Hybrid Imitation Learning: Combining Offline Trajectory-wise and Interactive State-wise Annotations
Practical deployments of imitation learning systems often learn simultaneously from offline and interactive feedback modalities kelly2019hg; hoque2021thriftydagger: for example, in autonomous driving zhang2017query; badue2021self; zhao2023real, the learner has access to some offline expert demonstrations to start with, and also receives interactive expert demonstration feedback in trajectory segments for subsequent finetuning. Motivated by this practice, we formulate the following problem:
@@
1:Input: MDP $\mathcal{M}$, state-wise expert annotation oracle $\mathcal{O}^{\mathrm{State}}$, Markovian policy class $\mathcal{B}$, online learning oracle $\mathbb{A}$, offline expert dataset $D_{\text{off}}$ of size $N_{\mathrm{off}}$, online budget $N_{\mathrm{int}}$
2: Initialize $\mathbb{A}$ with policy class $\mathcal{B}_{\text{bc}}:=\{\pi\in\mathcal{B}:\pi(s_{h})=a_{h},\ \forall h\in[H],\forall(s_{h},a_{h})_{h=1}^{H}\in D_{\text{off}}\}$.
@@ 4.1 Warm-Stagger: Algorithm and Analysis
We answer the above questions by proposing the Warm-Stagger algorithm, namely Algorithm 2. It extends Stagger to incorporate offline expert demonstrations, in that it constructs $\mathcal{B}_{\text{bc}}$, a restricted policy class that contains all policies in $\mathcal{B}$ consistent with all offline expert demonstrations (line 2). It subsequently performs online log-loss optimization on $\mathcal{B}_{\text{bc}}$ over state-action pairs collected online, where the state $s^{n}$ is obtained by rolling out $\pi^{n}$ in the MDP $\mathcal{M}$, and the action $a^{*,n}$ is annotated by the state-wise expert annotation oracle $\mathcal{O}^{\mathrm{State}}$. For analysis purposes, we introduce the following technical definition.
Definition 4 (Each-step policy completion).
Given a base policy class $\mathcal{B}$, define for each step $h\in[H]$
$\mathcal{B}_{h}:=\bigl\{\,\pi_{h}\;\big|\;\pi=(\pi_{1},\ldots,\pi_{H})\in\mathcal{B}\,\bigr\}.$
Then the each-step completion of $\mathcal{B}$ is defined as
$\tilde{\mathcal{B}}:=\bigl\{\,\pi=(\pi_{1},\ldots,\pi_{H})\;\big|\;\pi_{h}\in\mathcal{B}_{h}\text{ for all }h\in[H]\,\bigr\}.$
@@
If Warm-Stagger is run with a state-wise expert annotation oracle $\mathcal{O}^{\mathrm{State}}$, an MDP $\mathcal{M}$ where $(\mathcal{M},\pi^{\mathrm{E}})$ is $\mu$-recoverable, a policy class $\mathcal{B}$ such that deterministic realizability (Assumption 1) holds, and the online learning oracle $\mathbb{A}$ set as the exponential weight algorithm with each-step mixing policies $\bar{\pi}_{u}\in\bar{\Pi}_{\mathcal{B}}$, then it returns $\hat{\pi}$ such that, with probability at least $1-\delta$,
$J(\pi^{\mathrm{E}})-J(\hat{\pi})\leq O\left(\min\left(\frac{R\log(\tilde{B}/\delta)}{N_{\mathrm{off}}},\frac{\mu H\log(B_{\text{bc}}/\delta)}{N_{\mathrm{int}}}\right)\right),$
@@
Theorem 5 shows that Warm-Stagger finds a policy with suboptimality guarantee not signficantly worse than BC or Stagger: first, Behavior Cloning using the offline data only has a suboptimality of $O\left(\frac{RH\log(B/\delta)}{N_{\mathrm{off}}}\right)$ (cf. Theorem 2); second, Stagger without using offline data has a suboptimality of $O\left(\frac{\mu H\log(B/\delta)}{N_{\mathrm{int}}}\right)$ (cf. Theorem 3). We conjecture that the $\log\tilde{B}$ dependence may be sharpened to $\log B$; we leave this as an interesting open question.
@@
One may consider another baseline that naively switches between BC and Stagger based on a comparison between their bounds; however, such a baseline needs to know $R$ and $\mu$ ahead of time. In practice, we expect our Warm-Stagger to perform much better than this baseline, since it seamlessly incorporates both sources of data, and its design does not rely on theoretical bounds that may well be pessimistic.
@@ 4.2 On the Benefit of Hybrid Imitation Learning
Theorem 5 is perhaps best viewed as a fall-back guarantee for Warm-Stagger: its performance is not much worse than either of the baselines. In this section, we demonstrate that the benefit of hybrid feedback modalities can go beyond this: we construct an MDP motivated by practical challenges, in which hybrid imitation learning has a significantly better sample efficiency than both offline BC and interactive Stagger. Specifically, we prove the following theorem:
@@
For large enough $S,H$, there exists an MDP $\mathcal{M}$ with $S$ states, and expert policy $\pi^{\mathrm{E}}$ such that:
@@
Theorem 7 suggests that when $HS\gg\max(1,C)$, Warm-Stagger achieves expert-level performance with significantly lower cost than two baselines. To see this, observe that Warm-Stagger has a total cost of $O(S+C)$, which is much smaller than $\Omega(HS)$ by BC, and $\Omega(HSC)$ by Stagger.
@@
We now sketch our construction of MDP $\mathcal{M}$. $\mathcal{M}$ has an episode length $H\geq 50$, $H=\Omega(\log(S))$ and action space of cardinality greater than $10H$. For each state, one of the actions is taken by the expert; the rest are “wrong” actions. We illustrate $\mathcal{M}$’s state space on the left of Figure 2; specifically, it consists of the following subsets:
  • Expert ideal states $\mathbf{E}$, where $|\mathbf{E}|=N_{0}$: this can model for example, the agent driving stably on the edge of a cliff ross2011reduction, where any incorrect action transitions the agent to the unrecoverable absorbing state set $\mathbf{B}:=\{\mathbf{b}\}$ (e.g., falling off the cliff). Taking the expert action keeps the agent in $\mathbf{E}$ with high probability $(1-\beta)$, and with a small probability $\beta$, moves the agent to $\mathbf{E^{\prime}}$ (e.g., a safe slope).
  • Unrecoverable state $\mathbf{B}=\left\{\mathbf{b}\right\}$: a special absorbing state that is unrecoverable by any action (dead).
  • Expert recoverable states $\mathbf{E^{\prime}}$: this models the agent getting off from the edge of the cliff to a safe slope. When in $\mathbf{E^{\prime}}$, taking the expert action allows the agent to return to a uniformly sampled state in $\mathbf{E}$. Taking a wrong action from $\mathbf{E^{\prime}}$ leads to $\mathbf{B^{\prime}}:=\{\mathbf{b^{\prime}}\}$ (e.g., rest area).
@@
We now briefly justify each algorithm’s performance as stated in Theorem 7. (1) BC only observes expert actions in $\mathbf{E}$ and $\mathbf{E^{\prime}}$, but never in $\mathbf{b^{\prime}}$. As a result, near-expert performance at test time requires high coverage over $\mathbf{E^{\prime}}$; otherwise, BC’s trained policy will likely incur compounding errors and get trapped in $\mathbf{b}^{\prime}$. (2) Stagger suffers from a cold-start problem: early policies fail to explore $\mathbf{E}$ efficiently, and incorrect actions can cause transitions into $\mathbf{b}$. Consequently, coverage over $\mathbf{E}$ grows slowly, and the policy may still fail on unannotated states in $\mathbf{E}$ even with $\Omega(HS)$ queries. (3) Warm-Stagger benefits from offline data that fully covers $\mathbf{E}$, and uses a small number of interactions to visit $\mathbf{b^{\prime}}$ and query the expert, avoiding costly exploration in $\mathbf{E^{\prime}}$ while matching expert performance.
We also conduct a simulation of the aforementioned three algorithms in a variant of the above MDP with $N_{0}=200$, $N_{1}=1000$, $H=100$, and $\beta=0.08$, using another reward function that assigns a reward of 1 only when the agent visits the states in $\mathbf{E}$. Here, we let the online learning oracle $\mathbb{A}$ optimize 0-1 loss, which is equivalent to minimizing log loss under a deterministic learner policy class and discrete actions. Figure 2 shows return and state coverage as functions of the number of expert annotations, averaged over 200 runs.
We observe that: (1) BC exhibits slow improvement, as $\mathbf{b^{\prime}}$ remains unseen (and thus unannotated) throughout training, resulting in poor performance even with substantial coverage (e.g., $80\%$) over $\mathbf{E^{\prime}}$; (2) Stagger is sample-inefficient due to slow exploration over $\mathbf{E}$ states, consistent with the cold-start intuition; (3) Warm-Stagger (WS), when initialized with limited (200) offline (state, expert action) pairs, still needs to explore $\mathbf{E}$ first before it can safely reach $\mathbf{E^{\prime}}$ without failure; and (4) Warm-Stagger with sufficient offline coverage on $\mathbf{E}$ (e.g., initialized with 3200 offline (state, expert action) pairs) directly benefits from exploring $\mathbf{b^{\prime}}$ with immediate performance gain, and enables safe and even faster exploration than the expert in $\mathbf{E^{\prime}}$.
Figure 3: Sample and cost efficiency on MuJoCo tasks. The top row shows expected return vs. number of annotations ($C=1$); the bottom row shows performance under a cost-aware setting ($C=2$). Warm-Stagger (WS) is initialized with 1/8, 1/4, or 1/2 of the total annotation budget as offline demonstrations. Specifically, WS($n$) refers to WS with offline expert trajectory demonstrations of total length $n$. For a good range of $n$’s, WS($n$) matches Stagger in sample efficiency and outperforms the baselines when $C=2$.
@@ 5 Related Work
, pioneered in autonomous driving pomerleau1988alvinn, was solved by offline, state‑wise supervised learning in early works ross2010efficient; syed2010reduction and named Behavior Cloning (BC). A recent analysis by foster2024behavior employs trajectory-wise Hellinger distance to tighten the dependence of BC on the horizon at the trajectory level, although its sample complexity measured per state still grows quadratically with the horizon in the worst case. This shortcoming, often termed covariate shift or compounding error pomerleau1988alvinn, arises when imperfect imitation drives the learner to unseen states, resulting in a cascading sequence of mistakes. From a data collection perspective, this can be mitigated by noise‑injection approaches such as laskey2017dart; ke2021grasping. By leveraging additional environment interactions, generative‑adversarial IL methods ho2016generative; sun2019provably; ke2020imitation; spencer2021feedback frame learning as a two‑player game that aims to find a policy that matches expert’s state‑action visitation distributions. This setting is also known as “apprenticeship learning using inverse reinforcement learning” in earlier works  abbeel2004apprenticeship; syed2007game, which also starts from offline demonstrations and assumes the ability to interact with the environment MDP. Quantitative comparisons with these methods are beyond our scope, as they rely on extensive interactions with the MDP and access to a class of discriminator functions, while we focus on understanding the utility of state‑wise interactive annotations. This line work also include recent work of ren2024hybrid, who introduce “Hybrid Inverse Reinforcement Learning”, which leverages hybrid Reinforcement Learning to accelerate its inner loop of policy search; in contrast, our “hybrid” setting focus on utilizing heterogeneous data modalities. Recent offline imitation learning approaches chang2021mitigating; zeng2023demonstrations do not require MDP access but still require access to strong offline datasets, either with broad expert coverage or a large transition buffer. Our work assumes that interacting with the environment does not incur costs; we leave a detailed analysis that takes into account environment interaction cost as future work.
@@
, first proposed by ross2010efficient, allows the expert to provide corrective feedback to the learner’s action retroactively. Assuming low costs of expert recovery from mistakes, termed recoverability, DAgger ross2011reduction, and following works kim2013learning; ross2014reinforcement; sun2017deeply; cheng2018convergence; cheng2019predictor; rajaraman2021value outperform traditional BC both theoretically and empirically. However, this efficiency demands substantial annotation effort mandlekar2020human. Although DAgger ross2011reduction and some subsequent works sun2017deeply; rajaraman2021value; swamy2022minimax; foster2024behavior popularized the practice of annotating full trajectories, there has also been growing interest in state-wise annotations li2022efficient; sekhari2023selective; li2023agnostic, which appeared as early as ross2010efficient; judah2014active. In fact, practical applications of DAgger often adopt partial trajectory annotation in expert-in-the-loop mandlekar2020human; spencer2022expert; liu2022robot designs, as seen in zhang2016query; kelly2019hg; hoque2021thriftydagger; torok2025greedy, where issues such as inconsistencies caused by retroactive relabeling laskey2017comparing can be mitigated. These methods often leverage human- or machine-gated expert interventions to ensure safety during data collection zhang2017query; menda2017dropoutdagger, provide more targeted feedback menda2019ensembledagger; cui2019uncertainty, and enable learning on the fly shi2024yell. The use of selective state-wise queries aligns with our goal of promoting interactive imitation learning with efficient supervision and provable sample efficiency. We regard our contribution as providing a theoretical foundation for this increasingly popular paradigm of partial trajectory annotation.
@@
Many practical deployments of interactive learning systems do not start from tabula rasa; instead, prior knowledge of various forms is oftentimes available. Combining offline and interactive feedback has recently gained much popularity in applications such as instruction finetuning large language models dongrlhf; ouyang2022training, and bandit machine translation nguyen2017reinforcement. Many recent theoretical works in reinforcement learning try to quantify the computational and statistical benefit of combining offline and online feedback: for example, li2023reward; tanhybrid show provable reduction of sample complexity using hybrid reinforcement learning, using novel notions of partial coverage; songhybrid shows that under some structural assumptions on the MDP, hybrid RL can bypass computational barriers in online RL kane2022computational. Many works also quantify the benefit of utilizing additional offline data sources in the contextual bandit domain; for example, oetomo2023cutting; sharma2020warm; zhang2019warm study warm-starting contextual bandits using offline bandit data and supervised learning data. While some variants of DAgger zhang2017query; hoque2021thriftydagger also operate in a hybrid setting, our work focuses on a fundamental formulation that explicitly accounts for the cost asymmetry between offline and interactive annotations settles2008active, and, to the best of our knowledge, is the first to provide a rigorous framework with provable sample efficiency guarantees.

### v3

@@ 3.2 Experimental Comparison
We conduct a simple simulation study comparing the sample efficiency of log-loss Behavior Cloning Foster et al. (2024) and Stagger in four MuJoCo Todorov et al. (2012); Brockman et al. (2016) continuous control tasks with $H=1000$ and pretrained deterministic MLP experts Schulman et al. (2015, 2017). Considering MuJoCo’s low sensitivity to horizon length Foster et al. (2024), we reveal expert states one by one along consecutive trajectories for BC to allow fine-grained state-wise sample complexity comparison, while Stagger queries exactly one state per iteration by sampling from the latest policy’s rollout and updating immediately with the expert’s annotation. In Stagger, we implement the online learning oracle $\mathbb{A}$ so that it outputs a policy that approximately minimizes the log loss. In addition to log loss, we also include results with online learning oracle minimizing historical examples’ total square loss in Appendix G.2. We defer other implementation details to Appendix G.
@@
Figure 1: State-wise sample complexity comparison between Behavior Cloning and Stagger. Shaded areas show the 10th–90th percentile bootstrap confidence intervals DiCiccio and Efron (1996) over 10 runs. Stagger matches or exceeds BC with $50\%$ fewer annotations, achieving better state-wise annotation efficiency.
@@ 4 Hybrid Imitation Learning: Combining Offline Trajectory-wise and Interactive State-wise Annotations
Practical deployments of imitation learning systems often learn simultaneously from offline and interactive feedback modalities Kelly et al. (2019); Hoque et al. (2021): for example, in autonomous driving Zhang and Cho (2017); Badue et al. (2021); Zhao et al. (2023), the learner has access to some offline expert demonstrations to start with, and also receives interactive expert demonstration feedback in trajectory segments for subsequent finetuning. Motivated by this practice, we formulate the following problem:
@@
1:Input: MDP $\mathcal{M}$, state-wise expert annotation oracle $\mathcal{O}^{\mathrm{State}}$, stationary policy class $\mathcal{B}$, online learning oracle $\mathbb{A}$, offline expert dataset $D_{\text{off}}$ of size $N_{\mathrm{off}}$, online budget $N_{\mathrm{int}}$
2: Initialize $\mathbb{A}$ with policy class $\mathcal{B}_{\text{bc}}:=\{\pi\in\mathcal{B}:\pi(s_{h})=a_{h},\ \forall h\in[H],\forall(s,a)_{1:H}\in D_{\text{off}}\}$.
@@ 4.1 Warm-Stagger: Algorithm and Analysis
We answer the above questions by proposing the Warm-Stagger algorithm, namely Algorithm 2. It extends Stagger to incorporate offline expert demonstrations, in that it constructs $\mathcal{B}_{\text{bc}}$, a restricted policy class that contains all policies in $\mathcal{B}$ consistent with all offline expert demonstrations (line 2). It subsequently performs online log-loss optimization on $\mathcal{B}_{\text{bc}}$ over state-action pairs collected online, where the state $s^{n}$ is obtained by rolling out $\pi^{n}$ in the MDP $\mathcal{M}$, and the action $a^{*,n}$ is annotated by the state-wise expert annotation oracle $\mathcal{O}^{\mathrm{State}}$. For analysis, we introduce the following definitions.
Definition 4 (Non-stationary Markovian policies).
A non-stationary Markovian policy $\nu=(\nu_{1},\ldots,\nu_{H})$ is a collection of $H$ mappings, with each $\nu_{h}$ in $\Delta(\mathcal{A})^{{\mathcal{S}}}$, where upon rolling out $\nu$, at every step $h\in[H]$, the agent takes action $a_{h}$ by sampling from $\nu_{h}(\cdot\mid s_{h})$.
Definition 5 (Step-wise completion of stationary policy class).
For a stationary policy class $\mathcal{B}\subseteq\Delta(\mathcal{A})^{\mathcal{S}}$, define its step-wise completion to be a class of nonstationary Markovian policies:
$\tilde{\mathcal{B}}=\left\{\nu=(\nu_{1},\ldots,\nu_{H}):\nu_{h}\in\mathcal{B},\text{ for all }h\in[H]\right\}$
@@
If Warm-Stagger is run with a state-wise expert annotation oracle $\mathcal{O}^{\mathrm{State}}$, an MDP $\mathcal{M}$ where $(\mathcal{M},\pi^{\mathrm{E}})$ is $\mu$-recoverable, a policy class $\mathcal{B}$ such that deterministic realizability (Assumption 1) holds, and the online learning oracle $\mathbb{A}$ set as the exponential weight algorithm with each-step mixing policies, then it returns $\hat{\pi}$ such that, with probability at least $1-\delta$,
$J(\pi^{\mathrm{E}})-J(\hat{\pi})\leq O\left(\min\left(\frac{R\log(\tilde{B}/\delta)}{N_{\mathrm{off}}},\frac{\mu H\log(B_{\text{bc}}/\delta)}{N_{\mathrm{int}}}\right)\right),$ (3)
@@
Theorem 6 shows that Warm-Stagger finds a policy with suboptimality guarantee not signficantly worse than BC or Stagger: first, Behavior Cloning using the offline data has a suboptimality guarantee of $O\left(\frac{R\log(B/\delta)}{N_{\mathrm{off}}}\right)$ (cf. Theorem 2), and Warm-Stagger’s guarantee is worse by at most a factor of $H$; second, Stagger without using offline data has a suboptimality of $O\left(\frac{\mu H\log(B/\delta)}{N_{\mathrm{int}}}\right)$ (cf. Theorem 3), which is on par with the second term of Eq. (8). We conjecture that the $\log\tilde{B}$ dependence may be sharpened to $\log B$, and leave this as an interesting open question.
@@
One may consider another baseline that naively switches between BC and Stagger based on a comparison between their bounds; however, such a baseline needs to know $R$ and $\mu$ ahead of time. In practice, we expect Warm-Stagger to perform much better than this baseline, since it seamlessly incorporates both sources of data, and its design does not rely on theoretical bounds that may well be pessimistic.
@@ 4.2 On the Benefit of Hybrid Imitation Learning
Theorem 6 is perhaps best viewed as a fall-back guarantee for Warm-Stagger: its performance is not much worse than using either of the feedback source alone. In this section, we demonstrate that the benefit of hybrid feedback modalities can go beyond this: we construct an MDP, in which hybrid imitation learning has a significantly better sample efficiency than both offline BC and interactive Stagger. Specifically, we prove the following theorem:
@@
For large enough $S,H,A$, there exists an episodic MDP $\mathcal{M}$ with $S$ states, $A$ actions, and horizon $H$, and expert policy $\pi^{\mathrm{E}}$ such that:
@@
Theorem 8 suggests that when $HS\gg\max(1,C)$ and $C\gg\frac{1}{H}$, Warm-Stagger achieves expert-level performance with significantly lower cost than two baselines. To see this, observe that Warm-Stagger has a total cost of $O(S+C)$, which is much smaller than $\Omega(HS)$ by BC, and $\Omega(HSC)$ by Stagger.
@@
We now sketch our construction of MDP $\mathcal{M}$. $\mathcal{M}$ has an episode length $H=\Omega(\log(S))$ and action space of size greater than $10H$. For each state, one of the actions is taken by the expert; the rest are “wrong” actions. We illustrate $\mathcal{M}$’s state space on the left of Figure 2; specifically, it consists of the following subsets:
  • Unrecoverable state $\mathbf{B}:=\left\{\mathbf{b}\right\}$: a special absorbing state that is unrecoverable by any action (dead).
  • Expert ideal states $\mathbf{E}$, where $|\mathbf{E}|=N_{0}$: this can model for example, an autonomous driving agent driving stably on the edge of a cliff Ross et al. (2011), where any incorrect action transitions the agent to the unrecoverable state $\mathbf{b}$ (e.g., car falling off the cliff). Taking the expert action keeps the agent in $\mathbf{E}$ with high probability $(1-\beta)$, and with a small probability $\beta$, moves the agent to $\mathbf{E^{\prime}}$ (e.g., a safe slope).
  • Expert recoverable states $\mathbf{E^{\prime}}$: this models the agent driving outside the edge of the cliff in a safe slope. When in $\mathbf{E^{\prime}}$, taking the expert action allows the agent to return to a uniformly sampled state in $\mathbf{E}$. Taking a wrong action from $\mathbf{E^{\prime}}$ leads to reaching a recoverable state $\mathbf{b^{\prime}}$ (e.g., rest area).
@@
We now briefly justify each algorithm’s learning performance as stated in Theorem 8. First, BC only observes expert actions in $\mathbf{E}$ and $\mathbf{E^{\prime}}$, but never in $\mathbf{b^{\prime}}$. As a result, near-expert performance at test time requires high coverage over $\mathbf{E^{\prime}}$; otherwise, BC’s trained policy will likely incur compounding errors and get trapped in $\mathbf{b}^{\prime}$. Second, Stagger suffers from a cold-start problem: early policies fail to explore $\mathbf{E}$ efficiently, and incorrect actions can cause transitions into $\mathbf{b}$. Consequently, coverage over $\mathbf{E}$ grows slowly, and the policy may still fail on unannotated states in $\mathbf{E}$ even with $\Omega(HS)$ queries. Lastly, Warm-Stagger benefits from offline data that fully covers $\mathbf{E}$, and uses a small number of interactions to visit $\mathbf{b^{\prime}}$ and query the expert, avoiding costly exploration in $\mathbf{E^{\prime}}$ while matching expert performance.
We also conduct a simulation of the aforementioned three algorithms in a variant of the above MDP with $N_{0}=200$, $N_{1}=1000$, $H=100$, and $\beta=0.08$, using another reward function that assigns a reward of 1 only when the agent visits the states in $\mathbf{E}$. Here, we let the online learning oracle $\mathbb{A}$ optimize 0-1 loss on the data seen so far, which is equivalent to minimizing log loss under a class of deterministic policies and discrete actions. Figure 2 shows return and state coverage as functions of the number of expert annotations, averaged over 200 runs.
We observe that: (1) BC exhibits slow improvement, as $\mathbf{b^{\prime}}$ remains unseen (and thus unannotated) throughout training, resulting in poor performance even with substantial coverage (e.g., $80\%$) over $\mathbf{E^{\prime}}$; (2) Stagger is sample-inefficient due to slow exploration over states in $\mathbf{E}$, consistent with the cold-start intuition; (3) Warm-Stagger (WS), when initialized with limited (e.g., 200) offline (state, expert action) pairs, still needs to explore $\mathbf{E}$ first before it can safely reach $\mathbf{E^{\prime}}$ without failure; and (4) Warm-Stagger with sufficient offline coverage on $\mathbf{E}$ (e.g., initialized with 3200 offline (state, expert action) pairs) directly benefits from exploring $\mathbf{b^{\prime}}$ with immediate performance gain, and enables safe and even faster exploration than the expert in $\mathbf{E^{\prime}}$.
Figure 3: Sample and cost efficiency on MuJoCo tasks. The top row shows expected return vs. number of annotations ($C=1$); the bottom row shows performance in a cost-aware setting ($C=2$). Warm-Stagger (WS) is initialized with 1/8, 1/4, or 1/2 of the total annotation budget as offline demonstrations. Specifically, WS($n$) refers to WS with offline expert trajectory demonstrations of total length $n$. For a good range of $n$’s, WS($n$) matches Stagger in sample efficiency and outperforms the baselines when $C=2$.
@@ 5 Related Work
, pioneered in autonomous driving Pomerleau (1988), was solved by offline, state‑wise supervised learning in early works Ross and Bagnell (2010); Syed and Schapire (2010) and named Behavior Cloning (BC). A recent analysis by Foster et al. (2024) employs trajectory-wise Hellinger distance to tighten the dependence of BC on the horizon at the trajectory level, although its sample complexity measured per state still grows quadratically with the horizon in the worst case. This shortcoming, often termed covariate shift or compounding error Pomerleau (1988), arises when the learned policy’s imperfect imitation drives the learner to unseen states, resulting in a cascading sequence of mistakes. From a data collection perspective, this can be mitigated by noise‑injection approaches such as Laskey et al. (2017b); Ke et al. (2021). By leveraging additional environment interactions, generative‑adversarial IL methods Ho and Ermon (2016); Sun et al. (2019); Ke et al. (2020); Spencer et al. (2021) frame learning as a two‑player game, and aims to find a policy that matches expert’s state‑action visitation distributions. This setting is also known as “apprenticeship learning using inverse reinforcement learning” in earlier works Abbeel and Ng (2004); Syed and Schapire (2007). Quantitative comparisons with these methods are beyond our scope, as they rely on extensive interactions with the MDP and access to a class of discriminator functions, while we focus on understanding the utility of state‑wise interactive annotations. This line of works also include recent work of Ren et al. (2024), who introduce “Hybrid Inverse Reinforcement Learning”, which leverages hybrid reinforcement learning Song et al. to accelerate its inner loop of policy search; different from theirs, our “hybrid” setting focus on utilizing heterogeneous data modalities. Recent offline imitation learning approaches Chang et al. (2021); Zeng et al. (2023) do not require MDP access but still require access to offline datasets possibly collected by non-expert policies, either with broad expert coverage or a large transition buffer. Our work assumes that interacting with the environment does not incur costs; we leave a detailed analysis that takes into account environment interaction cost as future work.
@@
, first proposed by Ross and Bagnell (2010), allows the expert to provide corrective feedback to the learner’s action retroactively. Assuming low costs of expert recovery from mistakes ($\mu$-recoverability; recall Section 2), DAgger Ross et al. (2011), and following works Kim et al. (2013); Ross and Bagnell (2014); Sun et al. (2017); Cheng and Boots (2018); Cheng et al. (2019); Rajaraman et al. (2021) outperform traditional BC both theoretically and empirically. However, this efficiency demands substantial annotation effort Mandlekar et al. (2020). Although DAgger Ross et al. (2011) and some subsequent works Sun et al. (2017); Rajaraman et al. (2021); Swamy et al. (2022); Foster et al. (2024) popularized the practice of annotating full trajectories, there has also been growing interest in state-wise annotations Nguyen and Daumé III (2020); Li and Zhang (2022); Sekhari et al. (2023); Li and Zhang (2023), which appeared as early as Ross and Bagnell (2010); Judah et al. (2014). In fact, practical applications of DAgger often adopt partial trajectory annotation in expert-in-the-loop Mandlekar et al. (2020); Spencer et al. (2022); Liu et al. (2022) designs, as seen in Zhang and Cho (2016); Kelly et al. (2019); Hoque et al. (2021); Torok et al. (2025), where issues such as inconsistencies caused by retroactive relabeling Laskey et al. (2017a) can be mitigated. These methods often leverage human- or machine-gated expert interventions to ensure safety during data collection Zhang and Cho (2017); Menda et al. (2017), provide more targeted feedback Menda et al. (2019); Cui et al. (2019), and enable learning on the fly Shi et al. (2024). The use of selective state-wise queries aligns with our goal of promoting interactive imitation learning with efficient supervision and provable sample efficiency. We regard our contribution as providing a starting point for understanding this increasingly popular paradigm of partial trajectory annotations.
@@
Many practical deployments of interactive learning systems do not start from tabula rasa; instead, prior knowledge of various forms is oftentimes available. For example, combining offline data and interactive feedback has recently gained much popularity in applications such as training large language models to follow human instructions Dong et al. ; Ouyang et al. (2022), and bandit machine translation Nguyen et al. (2017). Many recent theoretical works in reinforcement learning try to quantify the computational and statistical benefit of combining offline and online feedback: for example, Li et al. (2023); Tan et al. show provable reduction of sample complexity using hybrid reinforcement learning, using novel notions of partial coverage; Song et al. shows that under some structural assumptions on the MDP, hybrid RL can bypass computational barriers in online RL Kane et al. (2022). Many works also quantify the benefit of utilizing additional offline data sources in the contextual bandit domain; for example, Oetomo et al. (2023); Sharma et al. (2020); Zhang et al. (2019) study warm-starting contextual bandit learning using offline bandit logged data or supervised learning examples. While some variants of DAgger Zhang and Cho (2017); Hoque et al. (2021) also operate in a hybrid setting, our work focuses on a fundamental formulation that explicitly accounts for the cost asymmetry between offline and interactive annotations Settles et al. (2008).